average score of 30student is 6 if 28 score 160 find the remaining of the 2 student
Answer by Cromlix(4381)

Average score of 30 students is 6.
Total score = 6 * 30 = 180
28 students score 160
180 - 160 = 20
This is what the two students scored.
Hope this helps.
:-)